The parallax of Reiltas has been measured to be 0.285 arcseconds.

  1. What is the distance to Reiltas in parsecs? in light-years?
  2. Suppose Reiltas' parallax was re-measured by the new Satellite for Parallax Accuracy Measurements (SPAM) to be 0.275 arcseconds. Is Reiltas closer or farther than previously determined? Explain your reasoning.
Answer #1

D(In parsec) = 1/(in arcsecond)

a) distance = 1/0.285 = 3.5087 parsec

in light year = 3.5087*3.26156 = 11.443 light year

b) at 0.275 arcseconds

D = 1/0.275 = 3.6363parsec =3.6363*3.26156 = 11.86 light year

so it is farther than previously determined distance